Assume that gold has one electron per atom to carry charge. The mass density of gold is 1.89×104 kg/m3 and its molecular weight is 197.0 g/mol. Calculate the drift speed of the electrons in a wire made of gold, with a radius of 0.57 mm, that carries a current of 1.36 A.
Answer #1

There is one electron per atom present in gold. Thus, gold is a monovalent atom The current density is defined as the ration
Answer #2

Drift velocity is,

v = l / q*A

where l = current

q = charge density

A = area = pi*r^2

Molar density = 1.89*10^7 / 197 = 95939.086 mol/m^3

Atomic density = 95939.086 * 6.02*10^23 = 5.77*10^28 atoms/m^3

Charge density q = 5.77*10^28*1.6*10^(-19) = 9.24*10^9 C/^3

Drift velocity v = 1.36 / 9.24*10^9*pi*(0.57*10^(-3))^2

v = 4.53*10^(-4) m/s

Answer